OpenStudy (anonymous):

Simplify. (4xy^2)^3(xy)^5

OpenStudy (anonymous):

https://www.mathway.com/

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\Huge (4xy^2)^3(xy)^5$$

OpenStudy (anonymous):

Thanks but I think I got it =)

OpenStudy (skullpatrol):

Did you use $$(ab)^m =a^mb^m$$

OpenStudy (skullpatrol):

Also $$(a^m)^n=a^{mn}$$

OpenStudy (skullpatrol):

and finally $$a^m*a^n=a^{m+n}$$